OpenStudy (anonymous):

Find the partial fraction decomposition. (-22)/(6x^2+5x-4)

ganeshie8 (ganeshie8):

do you know the method to decompose ?

ganeshie8 (ganeshie8):

@km313 ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

Okay. il walk you through.. :) you ready ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\(\huge \frac{-22}{6x^2+5x-4} \)

ganeshie8 (ganeshie8):

lets factor out the denominator

ganeshie8 (ganeshie8):

\(\huge = \frac{-22}{6x^2+8x - 3x-4}\) \(\huge = \frac{-22}{2x(3x+4) - 1(3x+4)}\) \(\huge = \frac{-22}{(3x+4)(2x-1)}\) so far ok ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

our goal is to write the given expression in the form as below : \(\huge = \frac{A}{(3x+4)} + \frac{B}{(2x-1)}\)

ganeshie8 (ganeshie8):

lets equate our expression to above

ganeshie8 (ganeshie8):

\(\huge \frac{-22}{(3x+4)(2x-1)} = \frac{A}{(3x+4)}+\frac{B}{(2x-1)}\)

ganeshie8 (ganeshie8):

@km313 you still wid me... ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\(\huge = \frac{-22}{(3x+4)(2x-1) } = \frac{A(2x-1) + B(3x+4)}{(3x+4)(2x-1)}\)

ganeshie8 (ganeshie8):

above simplification makes sense to u ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

\(\huge = \frac{-22}{(3x+4)(2x-1) } = \frac{2Ax-1 + 3Bx+4B}{(3x+4)(2x-1)}\)

ganeshie8 (ganeshie8):

\(\huge = \frac{-22}{(3x+4)(2x-1) } = \frac{x(2A+3B) + (4B-1)}{(3x+4)(2x-1)}\)

ganeshie8 (ganeshie8):

we just isolated the "x" term.. did u see it ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

lets compare "x" terms & constants 0 = 2A + 3B ---- (1) -22 = 4B-1 -----(2)

ganeshie8 (ganeshie8):

so far ok.. ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good :) solve (1) and (2) for A & B

ganeshie8 (ganeshie8):

can you do that and tell me the values of A & B ?

ganeshie8 (ganeshie8):

actually we did one mistake

ganeshie8 (ganeshie8):

we should get below equations 0 = 2A + 3B ---- (1) -22 = 4B-A -----(2)

ganeshie8 (ganeshie8):

solving above two equations, you should get : A = 6, B = -4 just plugin these values in our first equation, to get the final expression : \(\huge \frac{6}{(3x+4)} + \frac{-5}{(2x-1)}\)

ganeshie8 (ganeshie8):

this is the final decomposed partial expression. hope it helped you partially... .:))

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

welcome... (: